# Long-term Causal Inference via Modeling Sequential Latent Confounding

### v1

@@ 7 Conclusion
where the inequality $(a)$ is based on $(a+b)^{2}\leq 2(a^{2}+b^{2})$, and $(b)$ is based on Assumption 5.3. Under Assumption 5.3, $\mu_{S_{t}}^{O}$, $\mu_{S_{t}}^{E}$ and their estimates are all bounded, and thus $\hat{\omega}_{t}(\mathbf{x})$ and its estimates are also bounded, thus $(b)$ holds. Then, applying Lemma D.1, we know for a fixed time step $t$, $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}-\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}=O_{p}(n^{\frac{-\alpha}{2\alpha+d}}+n^{\frac{-\beta}{2\beta+d}})$. Then, for $\hat{r}(\mathbf{x})$, we have the effective sample is of size $(T-1)n$, and then combining with Eq. (30) and $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$, we obtain the desired result. ∎
Appendix E Proof of Theorem 5.5
@@
$\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})=O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$ (32)

### v2

@@ Appendix D Additional Experimental Details and Results
where the inequality $(a)$ is based on $(a+b)^{2}\leq 2(a^{2}+b^{2})$, and $(b)$ is based on Assumption 5.3. Under Assumption 5.3, $\mu_{S_{t}}^{O}$, $\mu_{S_{t}}^{E}$ and their estimates are all bounded, and thus $\omega_{t}(\mathbf{x})$ and its estimates are also bounded, thus $(b)$ holds. Then, applying Lemma H.1, we know for a fixed time step $t$, $\frac{\hat{\omega}_{t+1}(\mathbf{x})}{\hat{\omega}_{t}(\mathbf{x})}-\frac{\omega_{t+1}(\mathbf{x})}{\omega_{t}(\mathbf{x})}=O_{p}(n^{\frac{-\alpha}{2\alpha+d}}+n^{\frac{-\beta}{2\beta+d}})$. Then, for $\hat{r}(\mathbf{x})$, we have the effective sample is of size $(T-1)n$, and then combining with Eq. (30) and $\mathcal{R}_{n}^{*}(x)=O_{p}((\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}})$, we obtain the desired result. And attaining the oracle rate requires $(\frac{1}{(T-1)n})^{\frac{\eta}{2\eta+d}}\geq\max\{(\frac{1}{(T-1)n})^{\frac{\alpha}{2\alpha+d}},(\frac{1}{(T-1)n})^{\frac{\beta}{2\beta+d}}\}$, yielding $\min\{\alpha,\beta\}\geq\eta$. ∎
Appendix I Proof of Theorem 5.5
@@
$\hat{\tau}(\mathbf{x})-\tau(\mathbf{x})$ (32)
$=$ $O_{p}\left(n^{-\frac{\gamma}{2\gamma+d}}+n^{-\frac{\alpha}{2\alpha+d}}+n^{-\frac{\beta}{2\beta+d}}+\frac{\mu}{((T-1)n)^{\frac{\eta}{2\eta+d}}}+\frac{\mu}{((T-1)n)^{\frac{\alpha}{2\alpha+d}}}+\frac{\mu}{((T-1)n)^{\frac{\beta}{2\beta+d}}}\right).$